OpenStudy (anonymous):

Find the distance between the points (1,-6) and (4,-2)

OpenStudy (mathstudent55):

Heve you learned about the distance formula?

OpenStudy (anonymous):

d=√(x2-x1)^2+(y2-y1)^2

OpenStudy (anonymous):

that's the formula :)

OpenStudy (anonymous):

so just plug it in x2=4 x1=1 y2=-2 y1=1

OpenStudy (anonymous):

are you there marcy?

OpenStudy (anonymous):

yeah

OpenStudy (goldphenoix):

\[d=√(x2-x1)^2+(y2-y1)^2\] ^Just to make the formula neater. c:

OpenStudy (anonymous):

it's real simple keep it simple don't think to much into it. look at the formula I gave you and I broke down the points for you into x2 x1 y2 and y1

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

sorry i don't know how to make it neater i'm just trying to help i'll learn as i go :)

OpenStudy (anonymous):

your fine

OpenStudy (anonymous):

start with x2-x1 then y2-y1 keep them inside the ( )

OpenStudy (anonymous):

The answer is 10.90 or √119 using the distance formula.

OpenStudy (anonymous):

don't give away the answer she won't learn that way. :(

OpenStudy (anonymous):

is it 226

OpenStudy (anonymous):

d√(4-1)^2 + (-2-1)^2 you would get this

OpenStudy (anonymous):

y-y1=y2-y1/x2-x1(x-x1) =4x-3y-22=0

OpenStudy (anonymous):

good one

OpenStudy (anonymous):

edit the board and you could finish the problem stick to one set of the parentheses at a time

OpenStudy (mathstudent55):